A person notices a mild shock if the current along a path through the thumb and index finger exceeds 78 μA. Compare the maximum possible voltage without shock across the thumb and index finger with a dry-skin resistance of 3.5 ✕ 105 Ω and a wet-skin resistance of 1,950 Ω.

(a) the required resistance of the coil

(b) the length of wire you must use to wind the coil
Answer #1

a)

From ohms law ,

V=I*R

the maximum possible voltage without shock across the thumb = 78*10^-6 A*3.5*10^5 Ω

= 27.3 volts.

b)

the maximum possible voltage across index finger with a dry-skin resistance = 78*10^-6 A*1950 Ω

= 0.152 volts